# Self-Guidance: Boosting Flow and Diffusion Generation on Their Own

### v4

@@ Abstract
Proper guidance strategies are essential to achieve high-quality generation results without retraining diffusion and flow-based text-to-image models. Existing guidance either requires specific training or strong inductive biases of diffusion model networks, potentially limiting their applications. Motivated by the observation that artifact outliers can be detected by a significant decline in the density from a noisier to a cleaner noise level, we propose Self-Guidance (SG), which improves the image quality by suppressing the generation of low-quality samples. SG only relies on the sampling probabilities of its own diffusion model at different noise levels with no need of any guidance-specific training. This makes it flexible to be used in a plug-and-play manner with other sampling algorithms. We also introduce a more efficient approximation of SG, named SG-prev, which reuses the output from the immediately previous diffusion step to avoid doubling sampling time. We conduct experiments on text-to-image and text-to-video generation with different architectures, including UNet and transformer models. With open-sourced diffusion models such as Stable Diffusion 3.5 and FLUX, SG exceeds existing algorithms on multiple metrics, including both FID and Human Preference Score. SG-prev also achieves strong results over both the baseline and the SG with only one forward pass. Moreover, we find that SG and SG-prev both have a surprisingly positive effect on the generation of physiologically correct human body structures such as hands, faces, and arms, showing their ability of eliminating human body artifacts with minimal efforts. We will release our code along with this paper.
@@
1 Introduction
Figure 1: We train a diffusion model on a 1-dimensional toy example with data drawn from two separate modes at ±$\pm 1$. At different noise levels $t$, we fit and plot the distribution of generated samples in the reverse diffusion process. (a) The results on diffusion time $t=0.38$. The blue curve plots the distribution of generated samples at the current noise level $t$, while the green curve plots the distribution of samples generated at the noisier level $t+\delta(t)$. The red curve shows the obtained SG sampling distribution after applying the ratio of these two probabilities with $\omega=1$. (b) With various $\omega$, we show that artifact samples around the origin are suppressed more sharply with a larger value of $\omega$, while the density of samples around two groundtruth modes is boosted. (c) The groundtruth two-mode distribution at ±$\pm 1$.
@@ 1 Introduction
Specifically, for a diffusion model, we find that the reason for generating unnatural images is due to the inadequate denoising that could yield artifacts in the reverse diffusion process. As Figure 1 shows, let us consider a simple diffusion model trained on a 1-dimensional toy dataset with two separate modes at ±$\pm 1$ as the ground truth. At the noise level $t=0.38$ for example, the probability $p_{t}(x_{t})$ (i.e., the blue curve) is likely to draw artifact samples around the origin as it has a large density at $x_{t}=0$.
To relieve it, an intuitive idea is to take a look at the diffused probability $p_{t+\delta(t)}(x_{t})$ at a noisier level $t+\delta(t)$ (i.e., the green curve). We know that the samples at this level could be noisier and thus contain more severe artifacts. Therefore, by comparing between these two levels, one can assume that if the density at the cleaner level $t$ (although it is still high) declines significantly from that at the noisier level $t+\delta(t)$, the corresponding samples are likely to be artifacts, since the belief of sampling them as cleaner outputs has greatly dropped as reflected by the declining density. This is what we have seen in Figure 1(a)-(b), where the current sampling probability $p_{t}(x_{t})$ of the blue curve at $t=0.38$ declines from that of the green curve at the noisier level $t+0.2$ at the origin. This reveals that the samples around the origin could be artifacts, which is true. This decline can be measured by the ratio of these two probabilities, resulting in a new sampling strategy with
$p^{SG}_{t}(x_{t})\propto p_{t}(x_{t})\big{\{}\frac{p_{t}(x_{t})}{p_{t+\delta(t)}(x_{t})}\big{\}}^{\omega},$ (1.1)
@@
In this paper, we present such an inference-time sampling strategy called Self-Guidance (SG) since it only involves its own diffusion model at different noise levels. For image generation tasks, as shown in Fig. 2, compared with the benchmark diffusion sampling algorithm, applying Self-Guidance successfully removes unwanted artifacts in generated images, fixing errors on human fingers and other generation errors, eliminating irrelevant objects, and improving text-image consistency.
@@
In addition, SG is architecture-independent and can therefore be integrated into almost all existing diffusion and flow-based models regardless of varying model architectures. Unlike CFG [11] and AG [23], SG does not require additional guidance-specific training. SG is orthogonal to many other guidance approaches [11, 23, 1], and it can seamlessly work together with them in a plug-and-play manner for better performances.
@@
In Section 5.2, we apply SG and SG-prev on Flux.1-dev [29] and Stable-Diffusion 3.5 [52, 8], as well as a representative text-to-video model, CogVideoX[67]. We show that when using SG only, all models achieved improved FID [10] and higher human preference scores on the HPS v2.1[64] benchmark. Combining SG and SG-prev with other established guidance methods, including CFG and PAG, leading to new state-of-the-art diffusion generation results, achieving 20.74 on FID, 31.56 on HPSv2, and 5.7845 on Aesthetic Score. Experiments on text-to-video generation also confirm the solid performance improvements brought by SG.
@@ 2 Related Works
Despite improvements in overall image structure and class-label alignment, existing guidelines have a common problem: they still generate images with artifacts, like six-finger hands or twisted human bodies, which brings significant concerns in AI safety and user experiences. As a comparison, our proposed Self-Guidance significantly improves fine-grained details of generated images, completing weaknesses of existing guidance in an orthogonal way. We will introduce details of SG in Section 4.
3 Preliminary
@@
In this section, we introduce preliminary knowledge and notations about diffusion models [59, 13, 57]. The flow models [34, 32] share similar concepts that we put in the Appendix B. Assume we observe data from the underlying distribution $q_{d}(\bm{x})$. The goal of generative modeling is to train models to generate new samples $\bm{x}\sim q_{d}(\bm{x})$. The forward diffusion process of DM transforms $q_{0}=q_{d}$ towards some simple noise distribution,
$\mathrm{d}\bm{x}_{t}=\bm{F}(\bm{x}_{t},t)\mathrm{d}t+G(t)\mathrm{d}\bm{w}_{t},$ (3.1)
where $\bm{F}$ is a pre-defined drift function, $G(t)$ is a pre-defined scalar-value diffusion coefficient, and $\bm{w}_{t}$ denotes an independent Wiener process. A continuous-indexed score network $\bm{s}_{\varphi}(\bm{x},t)$ is employed to approximate marginal score functions of the forward diffusion process (3.1). The learning of score networks is achieved by minimizing a weighted denoising score matching objective [61, 59],
@@
$\|\bm{s}_{\varphi}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t}|\bm{x}_{0})\|_{2}^{2}\mathrm{d}t.$ (3.2)
Here the weighting function $\lambda(t)$ controls the importance of the learning at different time levels and $q_{t}(\bm{x}_{t}|\bm{x}_{0})$ denotes the conditional transition of the forward diffusion (3.1). After training, the score network $\bm{s}_{\varphi}(\bm{x}_{t},t)\approx\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$ is a good approximation of the marginal score function of the diffused data distribution. High-quality samples from a DM can be drawn by simulating generative SDE (3) which is implemented by replacing the score function $\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$ with the learned score network [59].
@@
Though the diffusion model has a solid theoretical interpretation, directly using its score functions to simulate the generative SDE often leads to sub-optimal performances, especially for class-conditioned generation. Assume $\bm{c}$ is a class label, and $\bm{s}_{\varphi}(\bm{x}_{t},t|\bm{c})$ is a conditional score network. The pioneering work Ho and Salimans [11] introduces the classifier-free guidance (CFG), which replaces the vanilla probability $q_{t}(\bm{x}_{t}|\bm{c})$ with a new one $\tilde{q}_{t}(\bm{x}_{t}|\bm{c})\coloneqq q_{t}(\bm{x}_{t})(\frac{q_{t}(\bm{x}_{t}|\bm{c})}{q_{t}(\bm{x}_{t})})^{\omega}$. $q_{t}(\bm{x}_{t})$ represents the unconditional distribution which can be implemented by inputting an empty label $\bm{\varnothing}$ as $q_{t}(\bm{x}_{t}|\bm{\varnothing})$. With this, the new score function turns to
$\tilde{\bm{s}}_{\varphi}(\bm{x}_{t},t|\bm{c})\coloneqq\bm{s}_{\varphi}(\bm{x}_{t},t|\bm{\varnothing})+\omega\big{\{}\bm{s}_{\varphi}(\bm{x}_{t},t|\bm{c})-\bm{s}_{\varphi}(\bm{x}_{t},t|\bm{\varnothing})\big{\}}$ (3.4)
@@
4 Self-Guidance
Figure 4: We train a flow-based diffusion model on a two-dimensional example with data sampled from a double-swirl pattern. We plot the distribution of the generated samples at the last four noise levels. The heat map in (a) shows the magnitude of predicted velocity without self-guidance, while the subfigures (b), (c), and (d) show those with Self-Guidance with various $\omega$ and noise levels. With a larger $\omega$, more artifact outliers are successfully removed. Meanwhile, a too large value of $\omega$ (e.g., $\omega=7$) could underweight (or in other words, oversuppress) the sampling from the original density $p_{t}(\bm{x}_{t}|c)$ that covers the swirl distribution. Thus, a suitable value of guidance scale is necessary to balance the sampling coverage and the artifact suppression.
@@ 4 Self-Guidance
Formally, when generating the image with a condition $c$, the self-guided generative distribution at noise level $t$ can be formulated as discussed in Section 1,
$p_{t}^{SG}(\bm{x}_{t}|\bm{c},t)\propto$ $\,p_{t}(\bm{x}_{t}|\bm{c})\bigg{\{}\frac{p_{t}(\bm{x}_{t}|\bm{c})}{p_{t+\delta(t)}(\bm{x}_{t}|\bm{c})}\bigg{\}}^{\omega}.$ (4.1)
This probability contains two parts. One is the original probability $p_{t}(x_{t}|c)$ that contains both the desired and artifact samples. The other is the ratio between probabilities of two diffusion times and it is used to suppress the artifacts. The guidance scale $\omega$ acts as the combination weight.
Figure 1 has already illustrated the influence of $\omega$ on the Self-Guidance sampling. Here we provide a more complex 2-dimensional example in Figure 4 to show the effects of Self-Guidance with various values of $\omega$. The groundtruth data are sampled from a double-swirl pattern. By comparing the diffusion samplings with and without Self-Guidance, it demonstrates the ability of SG to remove the artifact outliers that do not reside on the swirls. With an increasing value of $\omega$, the outlier artifacts are suppressed more sharply, which eventually leads to the high quality sampling of the groundtruth distribution. However, it is worth noting that when $\omega$ becomes too larger (e.g., $\omega=7$), the distribution of generated samples could start drifting away from the swirls. This is not surprising because too large $\omega$ could underweight the original density $p_{t}(\bm{x}_{t}|c)$ that aims to cover the samplings of the swirl data. This shows that a suitable guidance scale $\omega$ is necessary to balance between the sampling of true data density and the suppression of artifact outliers. In experiments, we will study the effect of its choices in the ablation.
Since diffusion models output the score instead of the diffused probability, one can transform Eq. 4.2 into the score function by the relation of $\bm{s}(\bm{x}_{t}|\bm{c},t)=\nabla_{x_{t}}\log p(\bm{x}_{t}|c)$. Formally, we start with the log-density ratio between two nearby timesteps:
$\log\frac{p_{t}(x_{t}|c)}{p_{t+\delta(t)}(x_{t}|c)}=\log p_{t}(x_{t}|c)-\log p_{t+\delta(t)}(x_{t}|c),$ (4.2)
and take the gradient with respect to $x$ to obtain the score difference:
@@
$+\omega\Big{\{}\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{s}(\bm{x}_{t}|\bm{c},t+\delta(t))\Big{\}},$ (4.4)
Fig. 3 illustrates how Self-Guidance is applied in practice at each denoising step during the reverse diffusion process. For a noisy latent $\bm{x}_{t}$ at timestep $t$, the model simultaneously predicts the noise at $t$ and at a noiser timestep $t+\delta(t)$. These two predictions are then linearly combined as in Eq. 4, and the combined prediction is subsequently used to denoise $\bm{x}_{t}$ and obtain the latent $\bm{x}_{t-1}$ for the next denoising step. This can be easily implemented by inference with the same model with two different timesteps within a batch. This means SG does not require an additional network or specific training tricks as in other guidance algorithms, but only uses the trained model itself. That is why we called the method Self-Guidance.
@@
                 $=d_{\text{cod}(\theta)}(\bm{x}_{t},\bm{c},t+\delta(t))$   if SG
@@
                         $+\omega_{CFG}*(\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{s}(\bm{x}_{t}|t))$
                         $+\omega_{PAG}*(\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{\widehat{s}}(\bm{x}_{t}|\bm{c},t))$
                         $+\omega_{SG}*(\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{s}(\bm{x}_{t}|\bm{c},t+\delta(t)))$
      $\bm{x}_{t-1}$ = scheduler.step($\bm{s}^{*}(\bm{x}_{t}|\bm{c},t),t,\bm{x}_{t}$)
@@
Algorithm 1 Diffusion model inference with Self-Guidance and other guidances
There are various choices of the shifted noise scale $\delta(t)$. In this paper, we consider a constant shift scale that is fixed independent of diffusion time $t$, and a dynamic one $\delta(t)=t/\sigma$ that becomes smaller as $t$ approaches 00. The latter allows a stronger Self-Guidance at the beginning of the diffusion sampling process. Moreover, we can also simply reuse the output distribution $\bm{s}(\bm{x}_{t+1}|\bm{c},t+1)$ from the previous diffusion step to approximate the SG guidance term, as in Eq. 4
$\bm{s}^{SG\_prev}(\bm{x}_{t}|\bm{c},t)\coloneqq$ $\,\bm{s}(\bm{x}_{t}|\bm{c},t)$
$+\omega\bigg{\{}\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{s}(\bm{x}_{t+1}|\bm{c},t+1)\bigg{\}}.$ (4.5)
This avoids doubling inference time as it does not require an additional forward pass at each diffusion step. We call this approximation method SG-prev. In experiments, we will compare these choices in both Text-to-Image and Text-to-Video models.
We detail how SG and SG-prev are used in inference in Algorithm 1, as well as if combined with other diffusion guidance methods, such as CFG and PAG. For a single step at the noise level $t$, the diffusion model applies different guidance individually by forwarding the model with the corresponding settings: unconditional prediction in CFG, perturbing the attention in PAG, a shifted timestep $t+\delta(t)$ in SG, and the previous timesteps output in SG-prev. These correct terms provide orthogonal improvements. Combined with their corresponding weight factors $\omega_{CFG},\omega_{PAG},\omega_{SG}$, the overall prediction is optimized to achieve higher consistency and control in the quality of the generation.
Discussion
Suppressing artifacts with $\frac{p_{t}(\bm{x}_{t}|\bm{c})}{p_{t+\delta(t)}(\bm{x}_{t}|\bm{c})}$ can be motivated from the heat equation. Specifically, under Gaussian smoothing with kernel $\mathcal{N}(0,\sigma^{2})$, the diffused density evolves as:
[…]
$=(p_{t}*\mathcal{N}(0,\sigma^{2}))(x).$ (4.6)
@@ Discussion
$\partial_{t}p_{t}(x)=\frac{1}{2}g(t)^{2}\Delta_{x}p_{t}(x),$ (4.7)
@@
$\Delta_{x}p_{t}(x_{\text{artifact}})>0.$ (4.8)
@@
$\partial_{t}p_{t}(x_{\text{artifact}})>0\quad\Rightarrow\quad p_{t+\delta(t)}(x_{\text{artifact}})>p_{t}(x_{\text{artifact}})$ (4.9)
@@
5 Experiments

### v5

@@ Abstract
Proper guidance strategies are essential to achieve high-quality generation results without retraining diffusion and flow-based text-to-image models. Existing guidance either requires specific training or strong inductive biases of diffusion model networks, which potentially limits their ability and application scope. Motivated by the observation that artifact outliers can be detected by a significant decline in the density from a noisier to a cleaner noise level, we propose Self-Guidance (SG), which can significantly improve the quality of the generated image by suppressing the generation of low-quality samples. The biggest difference from existing guidance is that SG only relies on the sampling score function of the original diffusion or flow model at different noise levels, with no need for any tricky and expensive guidance-specific training. This makes SG highly flexible to be used in a plug-and-play manner by any diffusion or flow models. We also introduce an efficient variant of SG, named SG-prev, which reuses the output from the immediately previous diffusion step to avoid additional forward passes of the diffusion network. We conduct extensive experiments on text-to-image and text-to-video generation with different architectures, including UNet and transformer models. With open-sourced diffusion models such as Stable Diffusion 3.5 and FLUX, SG exceeds existing algorithms on multiple metrics, including both FID and Human Preference Score. SG-prev also achieves strong results over both the baseline and the SG, with 50 percent more efficiency. Moreover, we find that SG and SG-prev both have a surprisingly positive effect on the generation of physiologically correct human body structures such as hands, faces, and arms, showing their ability to eliminate human body artifacts with minimal efforts. We have released our code at https://github.com/maple-research-lab/Self-Guidance.
@@
I Introduction
Figure 1: Radar chart of 5 image benchmarks on Flux.1. Background numbers indicate reference values corresponding to normalized radii (0.5, 0.75, 1.0)
Figure 2: We train a diffusion model on a 1-dimensional toy example with data drawn from two separate modes at $\pm 1$. At different noise levels $t$, we fit and plot the distribution of generated samples in the reverse diffusion process. (a) The results on diffusion time $t=0.38$. The blue curve plots the distribution of generated samples at the current noise level $t$, while the green curve plots the distribution of samples generated at the noisier level $t+\delta(t)$. The red curve shows the obtained SG sampling distribution after applying the ratio of these two probabilities with $\omega=1$. (b) With various $\omega$, we show that artifact samples around the origin are suppressed more sharply with a larger value of $\omega$, while the density of samples around two groundtruth modes is boosted. (c) The ground truth two-mode distribution at $\pm 1$.
@@ I Introduction
Specifically, for a diffusion model, we find that the reason for generating unnatural images is due to the inadequate denoising that could yield artifacts in the reverse diffusion process. As Figure 2 shows, let us consider a simple diffusion model trained on a 1-dimensional toy dataset with two separate modes at $\pm 1$ as the ground truth. At the noise level $t=0.38$ for example, the probability $p_{t}(x_{t})$ (i.e., the blue curve) is likely to draw artifact samples around the origin as it has a large density at $x_{t}=0$.
To relieve it, an intuitive idea is to take a look at the diffused probability $p_{t+\delta(t)}(x_{t})$ at a noisier level $t+\delta(t)$ (i.e., the green curve). We know the samples at this level could be noisier and thus contain more severe artifacts. Therefore, by comparing between these two levels, one can assume that if the density at the cleaner level $t$ (although it is still high) declines significantly from that at the noisier level $t+\delta(t)$, the corresponding samples are likely to be artifacts, since the belief of sampling them as cleaner outputs has greatly dropped as reflected by the declining density. This is what we have seen in Figure 2(a)-(b), where the current sampling probability $p_{t}(x_{t})$ of the blue curve at $t=0.38$ declines from that of the green curve at the noisier level $t+0.2$ at the origin. This reveals that the samples around the origin could be artifacts, which is true. This decline can be measured by the ratio of these two probabilities, resulting in a new sampling strategy
$p^{SG}_{t}(x_{t})\propto p_{t}(x_{t})\big\{\frac{p_{t}(x_{t})}{p_{t+\delta(t)}(x_{t})}\big\}^{\omega},$ (I.1)
@@
In this paper, we present such an inference-time sampling strategy called Self-Guidance (SG) since it only involves its own diffusion model at different noise levels. For image generation tasks, as shown in Fig. 3, compared with the benchmark diffusion sampling algorithm, applying SG successfully removes unwanted artifacts in generated images, fixing errors on human fingers and other generation errors, eliminating irrelevant objects, and improving text-image consistency.
@@
In addition, SG is architecture-independent and can therefore be integrated into almost all existing diffusion and flow-based models regardless of varying model architectures. Unlike CFG [11] and AG [23], SG does not require additional guidance-specific training. SG is orthogonal to many other guidance approaches [11, 23, 1], seamlessly works together with them in a plug-and-play manner for better performances.
@@
In Section V-B, we apply SG and SG-prev on Flux.1-dev [29] and Stable-Diffusion 3.5 [51, 8], as well as a representative text-to-video model, CogVideoX[66]. We show that when using SG only, all models achieved improved FID [10] and higher human preference scores on the HPS v2.1[63] benchmark. Combining SG and SG-prev with other established guidance methods, including CFG and PAG, leading to new state-of-the-art diffusion generation results, achieving 20.74 on FID, 31.56 on HPSv2, and 5.7845 on Aesthetic Score. Experiments on text-to-video generation also confirm the solid performance improvements brought by SG.
@@ II Related Works
Despite improvements in overall image structure and class-label alignment, existing guidelines have a common problem: they still generate images with artifacts, like six-finger hands or twisted human bodies, which brings significant concerns in AI safety and user experiences. As a comparison, our proposed Self-Guidance significantly improves fine-grained details of generated images, completing weaknesses of existing guidance in an orthogonal way. We will introduce details of SG in Section IV.
III Preliminary
@@
In this section, we introduce preliminary knowledge and notations about diffusion models [58, 13, 56]. The flow models [34, 32] share similar concepts that we put in Appendix -B. Assume we observe data from the underlying distribution $q_{d}(\bm{x})$. The goal of generative modeling is to train models to generate new samples $\bm{x}\sim q_{d}(\bm{x})$. The forward diffusion process of DM transforms $q_{0}=q_{d}$ towards some simple noise distribution,
$\mathrm{d}\bm{x}_{t}=\bm{F}(\bm{x}_{t},t)\mathrm{d}t+G(t)\mathrm{d}\bm{w}_{t},$ (III.1)
where $\bm{F}$ is a pre-defined drift function, $G(t)$ is a pre-defined scalar-value diffusion coefficient, and $\bm{w}_{t}$ denotes an independent Wiener process. A continuous-indexed score network $\bm{s}_{\varphi}(\bm{x},t)$ is employed to approximate marginal score functions of the forward diffusion process (III.1). The learning of score networks is achieved by minimizing a weighted denoising score matching objective [60, 58],
@@
$\|\bm{s}_{\varphi}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t}|\bm{x}_{0})\|_{2}^{2}\mathrm{d}t.$ (III.2)
Here, the weighting function $\lambda(t)$ controls the importance of the learning at different time levels and $q_{t}(\bm{x}_{t}|\bm{x}_{0})$ denotes the conditional transition of the forward diffusion (III.1). After training, the score network $\bm{s}_{\varphi}(\bm{x}_{t},t)\approx\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$ is a good approximation of the marginal score function of the diffused data distribution. High-quality samples from a DM can be drawn by simulating generative SDE (III) which is implemented by replacing the score function $\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$ with the learned score network [58].
@@
[…]
$\tilde{\bm{s}}_{\varphi}(\bm{x}_{t},t|\bm{c})\coloneqq\bm{s}_{\varphi}(\bm{x}_{t},t|\bm{\varnothing})+\omega\big\{\bm{s}_{\varphi}(\bm{x}_{t},t|\bm{c})-\bm{s}_{\varphi}(\bm{x}_{t},t|\bm{\varnothing})\big\}$ (III.4)
@@
IV Self-Guidance
Figure 5: Comparison of SG and SG-prev under the two-mode distribution. (a) - (e) are the results on different diffusion time from $t=0.75$ to $t=0.25$. The blue curve plots the distribution of generated samples at the current noise level $t$, the green curve plots the distribution of samples generated at the noisier level $t+\delta(t)$, while the gray curve plots the distribution of samples immediately generated at the immediately previous diffusion step. (f) The ground truth two-mode distribution at $\pm 1$.
Figure 6: We train a flow-based diffusion model on a two-dimensional example with data sampled from a double-swirl pattern. We plot the distribution of the generated samples at the last four noise levels. The heat map in (a) shows the magnitude of predicted velocity without self-guidance, while the subfigures (b), (c), and (d) show those with Self-Guidance with various $\omega$ and noise levels. With a larger $\omega$, more artifact outliers are successfully removed. Meanwhile, a too large value of $\omega$ (e.g., $\omega=7$) could underweight (or, in other words, oversuppress) the sampling from the original density $p_{t}(\bm{x}_{t}|c)$ that covers the swirl distribution. Thus, a suitable value of the guidance scale is necessary to balance the sampling coverage and the artifact suppression.
@@ IV Self-Guidance
Formally, when generating the image with a condition $c$, the self-guided generative distribution at noise level $t$ can be formulated as discussed in Section I,
$p_{t}^{SG}(\bm{x}_{t}|\bm{c},t)\propto$ $\,p_{t}(\bm{x}_{t}|\bm{c})\bigg\{\frac{p_{t}(\bm{x}_{t}|\bm{c})}{p_{t+\delta(t)}(\bm{x}_{t}|\bm{c})}\bigg\}^{\omega}.$ (IV.1)
This probability contains two parts. One is the original probability $p_{t}(x_{t}|c)$ that contains both the desired and artifact samples. The other is the ratio between probabilities of two diffusion times, and it is used to suppress the artifacts. The guidance scale $\omega$ acts as the combination weight.
Figure 2 has already illustrated the influence of $\omega$ on the Self-Guidance sampling. Here we provide a more complex 2-dimensional example in Figure 6 to show the effects of Self-Guidance with various values of $\omega$. The ground truth data are sampled from a double-swirl pattern. By comparing the diffusion samplings with and without Self-Guidance, it demonstrates the ability of SG to remove the artifact outliers that do not reside on the swirls. With an increasing value of $\omega$, the outlier artifacts are suppressed more sharply, which eventually leads to the high-quality sampling of the ground truth distribution. However, it is worth noting that when $\omega$ becomes too large (e.g., $\omega=7$), the distribution of generated samples could start drifting away from the swirls. This is not surprising because too large $\omega$ could underweight the original density $p_{t}(\bm{x}_{t}|c)$ that aims to cover the samplings of the swirl data. This shows that a suitable guidance scale $\omega$ is necessary to balance between the sampling of true data density and the suppression of artifact outliers. In experiments, we will study the effect of its choices in the ablation.
Since diffusion models output the score instead of the diffused probability, one can transform Eq. IV.2 into the score function by the relation of $\bm{s}(\bm{x}_{t}|\bm{c},t)=\nabla_{x_{t}}\log p(\bm{x}_{t}|c)$. Formally, we start with the log-density ratio between two nearby timesteps:
$\log\frac{p_{t}(x_{t}|c)}{p_{t+\delta(t)}(x_{t}|c)}=\log p_{t}(x_{t}|c)-\log p_{t+\delta(t)}(x_{t}|c),$ (IV.2)
and take the gradient to $x$ to obtain the score difference:
@@
$+\omega\Big\{\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{s}(\bm{x}_{t}|\bm{c},t+\delta(t))\Big\},$ (IV.4)
Fig. 4 illustrates how Self-Guidance is applied in practice at each denoising step during the reverse diffusion process. For a noisy latent $\bm{x}_{t}$ at timestep $t$, the model simultaneously predicts the noise at $t$ and at a noisier timestep $t+\delta(t)$. These two predictions are then linearly combined as in Eq. IV, and the combined prediction is subsequently used to denoise $\bm{x}_{t}$ and obtain the latent $\bm{x}_{t-1}$ for the next denoising step. This can be easily implemented by inference with the same model with two different timesteps within a batch. This means SG does not require an additional network or specific training tricks as in other guidance algorithms, but only uses the trained model itself. That is why we called the method Self-Guidance.
There are various choices of the shifted noise scale $\delta(t)$. In this paper, we consider a constant shift scale that is fixed independent of diffusion time $t$, and a dynamic one $\delta(t)=t/\sigma$ that becomes smaller as $t$ approaches 0. The latter allows a stronger Self-Guidance at the beginning of the diffusion sampling process. Moreover, we can also simply reuse the output distribution $\bm{s}(\bm{x}_{t+1}|\bm{c},t+1)$ from the previous diffusion step to approximate the SG guidance term, as in Eq. IV
$\bm{s}^{SG\_prev}(\bm{x}_{t}|\bm{c},t)\coloneqq$ $\,\bm{s}(\bm{x}_{t}|\bm{c},t)$
$+\omega\bigg\{\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{s}(\bm{x}_{t+1}|\bm{c},t+1)\bigg\}.$ (IV.5)
This avoids doubling inference time as it does not require an additional forward pass at each diffusion step. We call this approximation method SG-prev.
In Figure 5, we present a comparison between SG and SG-prev on the same one-dimensional toy example as in Figure 2. We find that the so-called SG-prev method is more efficient while bringing additional positive effects.
During image denoising, the early steps are mostly noise, and the image gradually becomes clear in the final steps. This is reflected by the probability density $p_{t}(x_{t})$ in Figure 5: at high-noise levels, the distribution is nearly Gaussian, so changes between steps are small; at low-noise levels, the distribution varies significantly, making the previous diffusion step probability $p_{t+1}(x_{t+1})$ (gray curve) near $x=0$ higher than $p_{t+\delta(t)}(x_{t})$ at the slightly noisier level $t+\delta(t)$ (green curve), which leads to stronger suppression of artifacts.
@@
         $d_{\text{cod}(\theta)}(\bm{x}_{t},\bm{c},t+\delta(t))$   if SG
@@
                    $\omega_{CFG}*(\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{s}(\bm{x}_{t}|t))$
                    $\omega_{PAG}*(\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{\widehat{s}}(\bm{x}_{t}|\bm{c},t))$
                    $\omega_{SG}*(\bm{s}(\bm{x}_{t}|\bm{c},t)-\bm{s}(\bm{x}_{t}|\bm{c},t+\delta(t)))$
[…]
@@
Algorithm 1 Diffusion model inference with Self-Guidance and other guidance
We detail how SG and SG-prev are used in inference in Algorithm 1, as well as if combined with other diffusion guidance methods, such as CFG and PAG. For a single step at the noise level $t$, the diffusion model applies different guidance individually by forwarding the model with the corresponding settings: unconditional prediction in CFG, perturbing the attention in PAG, a shifted timestep $t+\delta(t)$ in SG, and the previous timesteps’ output in SG-prev. These correct terms provide orthogonal improvements. Combined with their corresponding weight factors $\omega_{CFG},\omega_{PAG},\omega_{SG}$, the overall prediction is optimized to achieve higher consistency and control in the quality of the generation.
Some discussions about the motivation of SG
Besides the empirical motivation, Self-guidance also shares solid theoretical backgrounds. In this part, we show that suppressing artifacts with $\frac{p_{t}(\bm{x}_{t}|\bm{c})}{p_{t+\delta(t)}(\bm{x}_{t}|\bm{c})}$ can be motivated from the heat equation. Specifically, under Gaussian smoothing with kernel $\mathcal{N}(0,\sigma^{2})$, the diffused density evolves as:
$p_{t+\delta(t)}(x)$ $=\int p_{t}(y)\,\mathcal{N}(x;y,\sigma^{2}\delta(t))\,dy$
$=(p_{t}*\mathcal{N}(0,\sigma^{2}))(x).$ (IV.6)
@@
$\partial_{t}p_{t}(x)=\frac{1}{2}g(t)^{2}\Delta_{x}p_{t}(x),$ (IV.7)
@@
$\Delta_{x}p_{t}(x_{\text{artifact}})>0.$ (IV.8)
@@
$\partial_{t}p_{t}(x_{\text{artifact}})>0\quad\Rightarrow\quad p_{t+\delta(t)}(x_{\text{artifact}})>p_{t}(x_{\text{artifact}})$ (IV.9)
@@
V Experiments
TABLE I: Quantitative results of different Stable Diffusion models and Flux with various guidance. FID, Clip Score, and Aesthetic Score are evaluated on the MS-COCO 2017 validation subset. Bolded values highlight the best performance.